OpenStudy (anonymous):

FAN AND MEDAL!! Solve questions below with work!! 7 - what percent of 500 is 35? 8 - what percent of 84 is 147? 9 - 52 is what percent of 78? 10 - What percent of 126 is 42? 11 - 70 is what percent of 1,000? 12 - 84 is what percent of 252?

OpenStudy (mathstudent55):

Rewrite each question as a statement. Then translate into an equation and solve the equation. For "what percent", use x/100. "is" means "=" "of" means "*"

OpenStudy (anonymous):

ik but i also know there are 3 types of percents, type 1, 2, and 3

OpenStudy (mathstudent55):

Here is 7 solved as an example: 7 - what percent of 500 is 35? x percent of 500 is 35 x/100 * 500 = 35 5x = 35 x = 7 Answer: 7%

OpenStudy (mathstudent55):

Notice that all the problems above are the same type. In all of them, the question is "what percent".

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

wait @mathstudent55

OpenStudy (anonymous):

how would i do #9?

OpenStudy (mathstudent55):

9 - 52 is what percent of 78? Change question to statement and use x for unknown: 52 is x percent of 78 Now write an equation: 52 = x/100 * 78 52 = 78x/100 78x/100 = 52 78x/100 * 100/78 = 52 * 100/78 x = 5200/78 = 66 2/3 Answer: 66 2/3%

OpenStudy (anonymous):

ty!

OpenStudy (mathstudent55):

yw